OpenStudy (anonymous):

I am trying to find the differance in the problems... (-12)-2 and -12-2.why does one turn into a negative fraction and the other positive?

OpenStudy (cwrw238):

(-12)-2 = -12* -2 = 24 -12-2 = -14

OpenStudy (swissgirl):

\((-12)^{-2}=({-1 \over 12})*({-1 \over 12})={1 \over 24}\) \(-(12)^{-2}=-({1 \over 12})*({1 \over 12})=-{1 \over 24}\)

OpenStudy (swissgirl):

whooppsss 1/144 and -1/144

OpenStudy (anonymous):

thank you so much for your help